LINE MIXING IN ATMOSPHERIC OZONE

\underline{COREY CASTO} AND FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210-1106.

The modeling of the ozone spectrum and the measurement of relevant spectral line parameters is of significant interest for atmospheric spectroscopy. While line width and shift measurements are easily done at lower pressures, the phenomenon of line mixing can only be observed at atmospheric pressures. We model line mixing parameters as defined by the Rosenkranz method in the 160-270 GHz region using microwave ring-down spectra. These parameters, the challenge of measuring them, and the importance of using a global fit to exclude systematic error will be discussed.